OpenStudy (anonymous):

A=(1+r/n)nt p=3000 r=3% n=12 t=6 find a

jhonyy9 (jhonyy9):

r=0,03

jhonyy9 (jhonyy9):

A=(n+r)/n *(nt)=(n+r)*t=(12+0,03)*6=12,03*6=72,18